OpenStudy (anonymous):

x^2-y^2+6x+10y-17=0 A.) Write this equation in standard form. B.) Find Center, Verticies, foci, and slope of asymptotes. PLEASE HELP!!! AHHHHH

OpenStudy (amistre64):

(x+3)^2 - (y-5)^2 = 17-25+9 (x+3)^2 (y-5)^2 ------- - ------- = 1 1 1 center = (-3,5) foci = (-3+sqrt(2),5) (-3-sqrt(2),5) slope of asymps = 1 and -1

OpenStudy (anonymous):

Thanks so much you're a lfiesaver!!!

OpenStudy (amistre64):

verts = (-4,5) (-2,5)

OpenStudy (amistre64):

do we know if im right yet?

OpenStudy (anonymous):

no not yet...

OpenStudy (anonymous):

im sure you are though

OpenStudy (amistre64):

lol.... im pretty sure I am too, but ive been known to be wrong

OpenStudy (anonymous):

lol.... I'll trust you on this one... besides I know I can't do any better. Hyperbolas and elipses are my worst subjects

OpenStudy (anonymous):

I have more feel free to answer if you want!!!

OpenStudy (amistre64):

i can hold my own with them, but they do get tricky

OpenStudy (amistre64):

:) iin due time

OpenStudy (anonymous):

:)